OpenStudy (anonymous):

Please Help!! Medal Will Be Rewarded! Find an equation for the tangent line at the point (5, 1). x^2+5y^2-4x+2y=0

OpenStudy (anonymous):

To find the tangent line, you have to find the derivative of the equation. Can you find the derivative of x^2 + 5y^2 - 4x + 2y ?

OpenStudy (anonymous):

yes one moment

OpenStudy (anonymous):

well with implicit differentiation i get y'=-(x-2)/5y+1 is this correct

OpenStudy (anonymous):

or its 2x-4

OpenStudy (anonymous):

I didn't do Implict differentiation, and I got: I got 2x + 10y - 4 + 2 = 2x + 10y - 2 You may be more familier with this than me. Why would you use implicit differentiation.

OpenStudy (zzr0ck3r):

nono you are differentiating with respect to one variable

OpenStudy (zzr0ck3r):

you must use implicit

OpenStudy (anonymous):

Thats what i did

OpenStudy (anonymous):

did you see my answer? is it correct?

OpenStudy (anonymous):

Lol listen to zz

OpenStudy (zzr0ck3r):

x^2+5y^2-4x+2y=0 2x+10yy'-4+2y'=0 10yy'+2y'=-2x+4 y'(10y+2)=-2x+4 y'=(-2x+4)/(10y+2) y'=2/2(2-x)/(5y+1) = (2-x)/(5y+1)

OpenStudy (zzr0ck3r):

so yes you are right

OpenStudy (anonymous):

okay so would my answer be my final answer?

OpenStudy (zzr0ck3r):

now plug in the point to get slope, then use y=mx+b

OpenStudy (anonymous):

it would be -3/6 ?

OpenStudy (zzr0ck3r):

if you plug in (5,1) you get (2-5)/(5+1) -3/6=-1/2 so you have y=mx+b y = (-1/2)x+b now use the point again to find b 1=(-1/2)5+b 1=-5/2+b b = 1+5/2 = 7/2 so y = (-1/2)x+(7/2)

OpenStudy (anonymous):

oh, okay, I'm so into this extreme calculus anymore i forget the basics. Okay, thank you so much! totally understand this a lot more now!

OpenStudy (zzr0ck3r):

np, as long as you get it now:)